please help me with this problem:
angle w is equal to angle z where the measure of angle w is (11x-8) and the measure of angle z is (9x+4) and they are both acute angles
this problem is from chapter 6, lesson 2.6 and lesson 2.7 for geometry

Answer by elima(1433)

angle w is equal to angle z where the measure of angle w is (11x-8) and the measure of angle z is (9x+4) and they are both acute angles
w=11x-8
z=9x+4
w=z
11x-8=9x+4
11x-9x=4+8
2x=12
x=6
w=11(6)-8
w=66-8
w=58
z=9(6)+4
z=54+4
z=58
:)